OpenStudy (anonymous):

What is the value of ∫_0^1▒ln⁡x/(1+ln⁡x )^2 dx

OpenStudy (anonymous):

what is this computer language

OpenStudy (anonymous):

I want to integrate (log x)/(1 + log x) between the limits 0 to 1

OpenStudy (anonymous):

sorry (log x)/(1 + log x)^2

OpenStudy (anonymous):

\[\frac{\ln x}{(1+\ln x)^2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there is a singularity in \(x=\frac{1}{e}\)

OpenStudy (anonymous):

can we integrate it or not

OpenStudy (anonymous):

we can...im searchin for it...one of the users asked it before

OpenStudy (anonymous):

please tell me if u find it thanks

OpenStudy (anonymous):

i cant find it ...but note that\[\int \frac{\ln x}{(1+\ln x)^2} \text{d}x=\frac{x}{1+\ln x}+C\]and play around this\[\int_{0}^{1} \frac{\ln x}{(1+\ln x)^2} \text{d}x=\lim_{\epsilon \rightarrow 0} (\int_{0}^{\frac{1}{e}-\epsilon} \frac{\ln x}{(1+\ln x)^2} \text{d}x+\int_{\frac{1}{e}+\epsilon}^{1} \frac{\ln x}{(1+\ln x)^2} \text{d}x)\]

OpenStudy (anonymous):

@mukushla you going for Cauchy Principal value ? i was thinking By parts .

OpenStudy (anonymous):

thanks bt u give me direct answer can u explain it how it comes

OpenStudy (anonymous):

bt i think we cannot solve it by parts

OpenStudy (anonymous):

i think we can solve it by parts. Wolfram also says yes Parts can be used. http://www.wolframalpha.com/input/?i=integrate+%28ln%28x%29%2F%281%2Bln%28x%29%29%5E2

OpenStudy (anonymous):

thanks sami thank u very much